### v1

# Demand, and consumer surplus in the payday-loan market: Evidence from British Columbia

Abstract
This study examines how interest rate caps affect the demand for payday loans, using aggregate data from British Columbia (2012–2019) during which the province’s maximum fee was reduced from \$23 to \$17 and then to \$15 per \$100 borrowed. Estimating a linear demand function via OLS, we find that lowering interest rate caps significantly increases loan demand. We estimate that the \$8 decrease, from \$23 to \$15 per \$100, raised annual consumer surplus by roughly \$28.6 million (2012 CAD). A further reduction to \$14, starting in January 2025, would add another \$3.9 million per year. These results suggest that stricter interest rate caps can yield substantial consumer welfare gains.
JEL: D14, G21, L51.
Keywords: payday loans; interest rate cap; consumer surplus.
1 Introduction
Payday loans are typically extended to borrowers with limited liquidity. The high cost of this credit can lead to repeat borrowing and persistent debt cycles (Stegman (2007)). Users of these products in North America are disproportionately young and from marginalized communities; many have low or volatile incomes, precarious employment, and thin or impaired credit histories (Buckland (2018); Apgar and Herbert (2006); Visano (2008)). To reduce the risk of exploitative practices and protect vulnerable consumers, most jurisdictions in the United States and Canada have enacted payday lending regulations. One common measure is a cap on borrowing costs, aimed at curbing the harmful features of payday lending (Kaufman (2013); Dilay and Williams (2018)).
Payday lenders typically price at or near the prevailing cap (Robinson (2018); Fekrazad (2020)). For example, a 15% cap (i.e., a \$15 fee per \$100 borrowed for a two-week loan) translates into an annual percentage rate (APR) of approximately 391%. Where market power exists, price ceilings may enhance efficiency. In addition, interest rate caps can serve as a rudimentary form of social insurance by reallocating surplus from lower-marginal-utility lenders to higher-marginal-utility borrowers (Glaeser and Scheinkman (1998)).
However, direct evidence linking interest rates to the demand for payday loans remains limited. In this study, we exploit policy changes in British Columbia between 2012 and 2019 that reduced the maximum fee per \$100 borrowed from \$23 to \$17, and subsequently to \$15. Using provincial-level aggregate data on borrowing costs and total loan value, we estimate a linear demand curve for payday loans and calculate the associated welfare effects. We estimate that the \$8 reduction from \$23 to \$15 increased consumer surplus by about \$28.6 million per year (2012 CAD). A recent policy effective January 2025 further lowers the cap from \$15 to \$14 per \$100. We estimate that this will add roughly \$3.8 million (2012 CAD) in annual consumer surplus.
2 Data
The payday-loan industry in British Columbia is licensed and regulated by Consumer Protection BC, which compiles annual operational statistics from all licensed lenders. The available dataset includes each year’s average borrowing cost (the fee charged per \$100 loan on a two-week term), total loan volume in nominal dollars, and total loan volume in real 2012 CAD (adjusted for inflation). Our sample covers 2012 through 2023, and summary statistics are provided in Table 1 (loan volumes are rounded to the nearest \$0.1 million). Over this period, the regulated maximum fee per \$100 was progressively lowered from \$23 (2012–2016) to \$17 in 2017, and to \$15 from 2018 onward. In practice, lenders priced loans at or very near the cap in each year. The reported annual average borrowing cost can slightly exceed the nominal cap in some years due to reporting cycle misalignment.
Table 1: British Columbia payday-loan market statistics, 2012-2023
Year Borrowing cost Total loan volume Total loan volume
(per CAD 100) (nominal CAD) (2012 CAD)
2012 21.50 318.1 Mil 318.1 Mil
2013 21.70 351.4 Mil 344.4 Mil
2014 21.90 385.3 Mil 370.3 Mil
2015 21.70 340.9 Mil 321.2 Mil
2016 21.70 369.7 Mil 341.5 Mil
2017 19.30 397.3 Mil 359.8 Mil
2018 16.82 416.1 Mil 369.5 Mil
2019 15.30 441.5 Mil 384.4 Mil
2020 14.92 390.7 Mil 333.5 Mil
2021 14.69 271.5 Mil 227.2 Mil
2022 14.85 305.7 Mil 250.8 Mil
2023 14.92 327.2 Mil 263.2 Mil
Notes: Borrowing cost is the dollar fee per CAD 100 borrowed on 14-day terms. Loan volumes are annual totals and are rounded to the nearest 0.1 million CAD. Real dollars are deflated to 2012 CAD.
3 Model
3.1 Demand of payday loans
We assume that borrowers face a linear inverse demand curve for payday loans given by:
$Q=a-bp,\qquad b>0,$ (1)
where $p$ is the borrowing cost per CAD 100 (the fee per \$100 loan for a two-week term) and $Q$ is the total annual real loan volume (in year 2012 values). Under a binding interest rate cap, $p$ is effectively set by policy and $Q$ is determined by demand at that price.
3.2 Estimation of demand
To empirically estimate the demand curve, we focus on the period before 2020. We exclude observations from 2020–2023 because the Canadian government’s temporary COVID-19 income-support programs during those years materially altered consumers’ short-term credit needs. These developments could confound the price–quantity relationship in the later data. Given that we have only annual observations, adding additional control variables or policy dummies is not feasible (doing so would exhaust our degrees of freedom and obscure identification). We therefore restrict the sample to 2012–2019 and run a simple OLS regression of annual real loan volume (2012 CAD) on the actual borrowing cost per \$100.
Table 2: Demand estimation: real loan volume on borrowing cost (2012–2019)
DV: Real loan volume (2012 \$)
Borrowing cost (per \$100) -6,621,526**
(2,618,230)
Constant 483,509,000***
(52,700,000)
Observations 8
R-squared 0.5154
F-statistic 6.38
Prob >> F 0.0449
  • Notes: OLS regression of annual real loan volume (2012 CAD) on the borrowing cost per \$100, restricted to years with fee cap above the 2020 change (2012–2019; $N=8$). Robust standard errors in parentheses. $^{*}p<0.10$, $^{**}p<0.05$, $^{***}p<0.01$.
Although this yields only eight data points, prior research suggests that even a small sample can produce informative estimates when the underlying relationship is strong and stable (Jenkins and Quintana-Ascencio (2020)). Consistent with theory, the estimated slope coefficient is negative and statistically significant: $\hat{\beta}=-6.61$ million CAD (SE = 2.62 million). This implies that a \$1 increase in the fee per \$100 is associated with about a 6.61 million CAD decrease in annual loan volume. The estimated intercept is $\hat{\alpha}=483.5$ million CAD (SE = 52.7 million). These OLS results are reported in Table 2.
4 Consumer surplus gains from lower interest rate caps
We quantify consumer-surplus (CS) gains from two cap reductions: (i) lowering the cap from \$23 to \$15 per \$100 borrowed, and (ii) lowering it further from \$15 to \$14 per \$100 (implemented Canada-wide in January 2025). Using the estimated demand curve $\hat{Q}(p)=\hat{a}-\hat{b}p$ with $\hat{a}=483.509$ and $\hat{b}=6.621$ (where $Q$ is in millions of 2012 CAD and $p$ in dollars per \$100), predicted volumes are
$Q_{23}=\hat{Q}(23)=331.226~\text{million},\quad Q_{15}=\hat{Q}(15)=384.194~\text{million},\quad Q_{14}=\hat{Q}(14)=390.815~\text{million}.$
The CS change from reducing the cap from $p_{1}$ to $p_{2}<p_{1}$ equals the area under demand between $p_{1}$ and $p_{2}$:
$\Delta CS(p_{1}\to p_{2})=\frac{1}{100}\int_{p_{1}}^{p_{2}}\hat{Q}(p)\,dp=\frac{1}{100}\Big[\hat{a}(p_{1}-p_{2})-\frac{\hat{b}}{2}\big(p_{1}^{2}-p_{2}^{2}\big)\Big],$ (2)
where the factor $1/100$ converts the fee (quoted per \$100) to a per \$1 price for integration. Plugging in $\hat{a}$ and $\hat{b}$, we obtain $\Delta CS_{23\to 15}\approx 28.62$ million (2012 CAD) and $\Delta CS_{15\to 14}\approx 3.88$ million (2012 CAD) per year. The latter figure is the projected annual benefit from the new cap introduced in 2025.
5 Conclusion
Using a small set of annual aggregate data, we estimate a linear demand curve for payday loans and approximate the associated consumer surplus. Despite the coarse granularity of the data—and the necessarily approximate welfare calculations—the results are consistent and policy-relevant. In British Columbia, lowering the cap from \$23 to \$15 per \$100 borrowed coincided with a sizable increase in annual loan volume (from roughly \$331 million to \$384 million, 2012 CAD) and an estimated \$28.6 million gain in consumer surplus per year. A further cap reduction to \$14 per \$100 (implemented nationally in 2025) is projected to add about \$3.9 million annually. These findings suggest that tighter interest-rate ceilings can enhance borrower welfare without suppressing overall loan usage.
The contraction of storefront lenders alongside the rise of online providers points to heterogeneous cost structures: online firms may profitably extend credit at lower fees than brick-and-mortar lenders. Nevertheless, policymakers should account for supply-side adaptation when considering further cap reductions. Much more can be learned from future research using microdata to examine heterogeneity across borrower segments, default risk, and lender margins, and to assess substitution toward other forms of credit. Overall, our evidence supports carefully calibrated caps as a viable consumer-protection tool in high-cost credit markets—provided regulators continue to monitor market composition and credit access as the industry evolves.